One of your friends purchased a zero coupon corporate bond (i.e., a bond that has no interest payments) for $4,850. The bond has a face value of $50,000 and is due in 15 years. If the bond is held to maturity, what rate of return will your friend make on the investment?

Question 22 options:

11.55%

The rate of return cannot be calculated with the information given.

16.83%

931.00%
Answer #1

Option C. 16.83%

The formula to calculate the,

The Current value of the zero coupon bond = Face value / (1+rate of return)^time period

Now, putting the values in the expression

4850 = 50000/(1+rate of return)^15

(1+rate of return)^15 = (50000/4850)

1+rate of return = (50000/4850)^(1/15)

rate of return = (50000/4850)^(1/15) - 1

= 0.168284

Therefore, rate of return = 16.83% approx.

Please don't forget to rate the answer if its helpful, thank you.